OpenStudy (anonymous):

Simplify the following expression. 4x − 6y + 2z − 5x − 3y + 8z

jimthompson5910 (jim_thompson5910):

4x - 5x = ???

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

more like -1x or just -x

jimthompson5910 (jim_thompson5910):

-6y - 3y = ???

OpenStudy (anonymous):

-9y

OpenStudy (anonymous):

-x-9y+10z=0

jimthompson5910 (jim_thompson5910):

and 2z +8z = ???

OpenStudy (anonymous):

10z

jimthompson5910 (jim_thompson5910):

So 4x − 6y + 2z − 5x − 3y + 8z simplifies to -x - 9y + 10z

OpenStudy (anonymous):

thank u

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

2x − 8y + 3x2 + 7y − 12x